# Phantom matter: a challenging solution to the cosmological tensions

### v1

@@ 5 Discussion and conclusions
As indicated, bubbles of PM are connected with the process of attaining the de Sitter (dS) era both in the early and the late universe. As noted above, during the transit from the provisional phantom vacuum into the true and stable (dS) vacuum, the PM bubbles being tunneled in the late universe are characterized by positive pressure and hence offer a fertile field for the growing of unsuspected structures. This mechanism, therefore, provides a tantalizing theoretical framework capable of explaining physically the appearance of a late time AdS epoch (or epochs) before the usual dS epoch is eventually attained  [^3] This is in contradistinction to the alternative $\Lambda_{s}$ scenario analyzed in  [39], although theoretical attempts at justifying it have appeared recently [96, 97, 98].. In point of fact, we do not expect that these PM bubbles are perfect AdS phases and for this reason we have left the EoS $w_{X}$ as a free parameter, whose fitted value is $w_{X}\simeq-1.16$ (cf. Table 1). Since the bubble is transitory, the subsequent DE phase below $z_{t}$ carries another EoS which we fit to be quintessence-like: $w_{Y}\simeq-0.90$ (see Table 1). We have performed a comparative fit of the $w$XCDM and $\Lambda_{s}$ scenarios using the same data sets and have found a significantly better fit for the PM scenario, as shown in Table 1.
@@
As remarked, the PM phase (or phases) with negative energy density and positive pressure are left behind during the cosmic evolution at relatively large redshifts of order $1-10$, whereas the subsequent evolution (closer to our time) flips into the quintessence regime. Although we have used an abrupt $\theta$-function behavior to connect the two EoS regimes, in practice the process is continuous since the Chern-Simons condensates eventually dominate and restore the normal vacuum phase with positive energy. To better understand how the PM bubbles with positive pressure enhance the formation of large (unsuspected) structures in the relatively distant the past, it is useful to analyze the equation for the matter density contrast, $\delta_{m}=\delta\rho_{m}/\rho_{m}$ in the presence of PM. We shall not provide details of the perturbations equations here (see [51]), but we wish at least to highlight this important aspect associated to the presence of PM in a more quantitative way. Before the transition at $z_{t}$ (i.e. at $z>z_{t}$), the equation for $\delta_{m}$ reads
with the primes denoting derivatives with respect to the scale factor. PM has negative energy density ($\Omega_{X}<0$) and positive pressure (due to $w_{X}<-1<0$) and, therefore, it induces a decrease of the friction term and an increase of the Poisson term in Eq. (4). Both effects are in harmony and cause, given some fixed initial conditions, an enhancement of the structure formation processes in the PM bubbles. Notice that this would not occur for ordinary phantom DE (cf. Fig. 1, for which the friction term gets enhanced and the Poisson term suppressed, i.e. just opposite to PM. Thus, quintessence suppresses LSS formation.
Moreover, because the energy of the $X$ entity is negative (and non-negligible at high redshift, in absolute value), this enforces a higher value of the expansion rate $H$ in the quintessence stage in order to preserve the angular diameter distance to the last scattering surface, which is essentially fixed by the precise measurement of $\theta_{*}$ (the angular size of sound horizon) by Planck [55] and the standard physics before recombination [37]. This explains why $H_{0}$ is found larger than in the $\Lambda$CDM in our PM scenario. Quantitatively, our value of $H_{0}$ (cf. Table 1) is in full agreement with that of SH0ES ($H_{0}=73.04\pm 1.04$ km/s/Mpc)[57] to within only $\sim 0.25\sigma$. The Hubble tension is therefore completely washed out, also when it is formulated in terms of the absolute magnitude of SNIa: our fitted value of $M$ and that of SH0ES ($M=19.253\pm 0.027$ mag) differ by only $\sim 0.6\sigma$. At the same time the rate of LSS formation becomes suppressed below $z_{t}\sim 1.5$ during the quintessence-like regime, as explained above, which is in accordance with the observations. In Table 1 and, more graphically, in Fig. 2, we can see that the amplitude of the power spectrum at linear scales that is preferred by the data is pretty similar in all the models under study. Indeed, in all cases we find values of $\sigma_{12}\sim 0.78$. However, at a finer level, Table 2 in the Appendix tells us that the $w$XCDM is able to describe better the LSS data than $\Lambda$CDM, what means that this model is able to produce lower values of $f(z)\sigma_{12}(z)$ and, hence, of $f(z)$ since $R_{12}=12$ Mpc stays fixed. This can again be understood by looking at the equation for the density contrast, whose form at $z<z_{t}$ is identical to Eq. (4), but with the replacements $\Omega_{X}\to\Omega_{Y}$ and $w_{X}\to w_{Y}$,
@@
It is clear from Fig. 2 that the values of the matter density parameter $\Omega_{m}^{0}=\Omega_{m}(a=1)$ in the $w$XCDM are a way smaller than in $\Lambda$CDM. This obviously translates into larger values of $\Omega_{Y}(a)>0$ (higher than $\Omega_{\Lambda}^{0}$ in the $\Lambda$CDM), which makes matter fluctuations to grow less efficiently in our model after the transition (at $z<z_{t}$), obviously a welcome feature. As previously mentioned, despite the presence of PM in the past, the value of $\sigma_{12}$ in $w$XCDM is not significantly different ($\sim 0.4\sigma$) from that of $\Lambda$CDM. This is due to the slower increase of matter fluctuations in the late-time universe and the smaller value of the amplitude of primordial fluctuations found in $w$XCDM, which somehow compensate the previous effects, cf. again Table 1 and Fig. 4.
@@
From our numerical analysis we can see that both models $w$XCDM and $\Lambda_{s}$ offer a dramatic reduction of the cosmological tensions. This is quantitatively apparent from Table 1. One could say that the Hubble and growth tensions virtually disappear in these models. However, there are also quantitative (and conceptual) differences between them. In terms of the information criteria, we find that the two models $w$XCDM and $\Lambda_{s}$ are very strongly preferred over $\Lambda$CDM. Numerically, however, the $w$XCDM solution provides a significantly better global fit. For this model we find $\Delta$DIC,$\Delta$AIC55greater-than-or-equivalent-toabsent55$\gtrsim$ 55≳ 55, whereas for $\Lambda_{s}$ we obtain $\Delta$DIC,$\Delta$AIC40greater-than-or-equivalent-toabsent40$\gtrsim$ 40≳ 40. The difference of $15$ units is quite substantive and indicates (in the common parlance of the information criteria) that there is very strong evidence that model $w$XCDM is preferred to model $\Lambda_{s}$ (for the same data). Worth noticing is also the fact that our own results for $\Lambda_{s}$ agree with those originally reported in [39], despite the existing differences in the data sets employed in the two analyses. Apart from our using data on CCH, which were not considered in [39], we employ data on $f(z)\sigma_{12}(z)$ instead of the weak lensing (WL) data from KiDS. We do not use WL data in our analysis in order not to compromise the computation of the non-linear matter power spectrum, which might not be sufficiently well under control in the models under study. Our fitting results, though, are completely consistent with theirs, leading also to similar values of $S_{8}\sim 0.78$ (as we have checked), the latter being a more natural LSS parameter in the context of WL analyses, but not in our approach.
@@
Perhaps the most remarkable outcome of our analysis is that in the light of the overall fit quality of our results leading to such an outstanding support to the composite DE models from the information criteria, it should be fair to conclude that the standard $\Lambda$CDM model appears to be comparatively ruled out at a very high confidence level. Let us analyze the degree of impact from each data source on the previous statement. It can be inferred from Table 2 in the Appendix, where a detailed breakdown of the different contributions is reckoned. The angular/transversal BAO data certainly contributes in a significant way, as could be expected from the analysis of[37]. However, it is by no means the leading contribution, as it is only responsible for roughly $\sim 18\%$ of the success. A close inspection of the mentioned table shows that the composite DE models beat actually the $\Lambda$CDM in every single observable, not only in those driving the tensions. Particularly noticeable is the sizeable impact from such a solid asset of basic data as the CMB and SNIa+SH0ES observations, which are responsible for about $70\%$ of our successful fitting result, whereas the influence from LSS data is less than $10\%$; and, finally, that of CCH is marginal (only a few percent). The bulk of the successful fit therefore hinges upon the most fundamental cosmological data, which is truly remarkable. The $\Lambda$CDM is not only unable to explain the large values of $H_{0}$ measured by SH0ES or to further slow down the structure formation rate at low redshift, but it also introduces other tensions, e.g. with the value of the (reduced) cosmological matter parameter $\omega_{\rm m}=\omega_{b}+\omega_{\rm dm}$ that is preferred by the Planck data in models which assume standard model physics before the decoupling of the CMB photons. These values can be read off from Table 1, namely $\omega_{m}=0.142\pm 0.001$ for the $w$XCDM and $\omega_{m}=0.138\pm 0.001$ for the standard $\Lambda$CDM. These values differ by $2.8\sigma$.
@@
Figure 3: Confidence regions in the $w_{Y}-w_{X}$ plane of the $w$XCDM model, and the corresponding one-dimensional posterior distributions. The dotted lines are set at $w_{Y}=-1$ and $w_{X}=-1$. The intersection of the horizontal and vertical lines in the contour plot corresponds to the $\Lambda_{s}$ model, which falls $\sim 3\sigma$ away from the preferred region of the $w$XCDM. See the main text for further comments.
@@
Focusing now on the $w$XCDM parameters, in Fig. 3 we show the constraints obtained in the EoS plane $w_{Y}$-$w_{X}$, which involves two of the characteristic new parameters of the $w$XCDM model (the third one being $z_{t}$, shared with $\Lambda_{s}$). The central value of $w_{X}=-1.16$ falls in the phantom region (in fact, PM region since $\Omega_{X}<0$) but is compatible with $-1$ at $\sim 1\sigma$ CL. There is, in contrast, a non-negligible ($\sim 3\sigma$) preference for a quintessence-like evolution of the DE for the low redshift range nearer to our time $z<z_{t}$: $w_{Y}=-0.90\pm 0.03$ (see also Fig. 3). Our preference for quintessence-like behavior in this region can be contrasted with the situation in the $\Lambda_{s}$, where a simple $\Lambda$-term is assumed below the transition redshift (also above that redshift, with $\Lambda\to-\Lambda$). Now the fact that our fit with model $w$XCDM proves significantly better than with model $\Lambda_{s}$ (as commented above) suggests that indeed the quintessence option in the low redshift range is strongly preferred. In addition, our result aligns perfectly well with the one obtained from the analysis of the Pantheon+ data in the context of the flat $w$CDM parametrization [56], which showed that SNIa data alone lead to an EoS $w=-0.89\pm 0.13$ after marginalizing over $\Omega_{m}^{0}$. This marginalized result, though, is still compatible with a cosmological constant ($w=-1$). To explain why in our case is more focused on the quintessence domain, let us note that for $w$XCDM we find a tight constraint on the matter density parameter, strongly favoring the region of small values, $\Omega_{m}^{0}=0.269\pm 0.005$. This is induced by two facts: (i) CMB prefers values of $\omega_{m}\sim 0.142$ in models with standard pre-recombination physics; and (ii) the large values of $H_{0}\sim 73$ km/s/Mpc measured by SH0ES. Combining the aforementioned tight constraint on $\Omega_{m}^{0}$ with the constraints in the $w-\Omega_{m}^{0}$ plane obtained for the $w$CDM from the analysis of the Pantheon+ data alone (cf. the contours in cyan of Fig. 9 of [56]) we can break the existing degeneracy in that plane and explain the $\sim 3\sigma$ evidence for quintessence DE that we find in $w$XCDM for the low redshift range. For $w$XCDM such low values of $\Omega_{m}^{0}$ are not favored by Planck, since they are in tension with the angular diameter distance to the last scattering surface, see again Fig. 9 of [56]. Remarkably, this problem is fully averted in the $w$XCDM model owing precisely to the presence of PM in the high redshift stretch $z>z_{t}$, cf. Fig. 4, which compensates for the decrease of the contribution to the angular diameter distance in the low redshift domain. On the other hand, in the $\Lambda_{s}$ case we also find similar low values of $\Omega_{m}^{0}$, but to connect our $w$XCDM with this model we have to set $w_{Y}$ to $-1$ (rigid cosmological constant), so the model falls outside the $3\sigma$ region obtained with $w$CDM using SNIa data alone [56]. This issue is also solved in the $w$XCDM and explains the improvement found with respect to the $\Lambda_{s}$ model. The full 5-year data set of high-redshift SNIa from the Dark Energy Survey will most likely increase the aforesaid signal, since they also find constraints in the region of interest, $(\Omega_{m}^{0},w)=(0.264^{+0.074}_{-0.096},-0.80^{+0.14}_{-0.16})$ [99]. Hence, it will be of utmost importance to study the impact of this data set once it becomes publicly available.

### v3

@@ 5 Discussion
As we have mentioned, bubbles of PM are connected with the process of attaining the de Sitter (dS) era both in the early and in the late universe. During the transit from the provisional phantom vacuum into the true and stable (dS) vacuum, the PM bubbles being tunneled in the late universe are characterized by positive pressure and hence offer a fertile field for the growing of unsuspected structures. This mechanism, therefore, provides a tantalizing theoretical framework capable of explaining physically the appearance of a late time AdS epoch (or epochs) before the usual dS epoch is eventually attained  [^5] This is in contradistinction to the alternative $\Lambda_{s}$ scenario analyzed in (Akarsu et al., 2023), although theoretical attempts at justifying it have appeared recently (Anchordoqui et al., 2024a; Akarsu et al., 2024b, a; Anchordoqui et al., 2024b).. In point of fact, we do not expect that these PM bubbles are perfect AdS phases (nor that the subsequent stage is necessarily characterized by a rigid $\Lambda>0$) and for this reason we have left the EoS parameter $w_{X}$ as a free parameter, whose fitted value turns out to be $w_{X}\simeq-1.16$ (cf. Table 1), although with $\rho_{X}<0$ (in contrast to conventional phantom DE). Since the bubble is transitory, the ensuing DE phase below $z_{t}$ carries another EoS which we fit it to be quintessence-like: $w_{Y}\simeq-0.90$ (see Table 1). We have performed a comparative fitting analysis of the $w$XCDM and $\Lambda_{s}$ scenarios using the same data sets and have found a significantly better fit quality for the PM scenario, as shown in Table 1. We shall come back to these results in a moment.
The PM phase (or phases) with negative energy density and positive pressure are left behind during the cosmic evolution at relatively large redshifts of order $1-10$, and in these places they can leave a sort of oasis rich of large scale structures, whereas the consecutive evolution (closer to our time) flips into the quintessence regime. Although we have used an abrupt $\theta$-function behavior to connect the two EoS regimes, in actual fact the process is continuous since the Chern-Simons condensates eventually dominate and restore the normal vacuum phase with positive energy. To better understand, in a quantitative way, how the PM bubbles with positive pressure may enhance the formation of (unsuspected) large scale structures in the relatively distant past, it is useful to analyze the differential equation for the matter density contrast in the presence of PM. This is a key aspect of the role played by PM for potentially helping to solve the cosmological tensions. Before the transition at $z_{t}$ (i.e. for $z>z_{t}$), the equation for $\delta_{m}$ reads [^6] As indicated, we shall not provide computational details here. For a full fledged exposition, including the complete set of coupled perturbations equations of the $w$XCDM model, see (Gómez-Valent & Solà Peracaula, 2024).
with the primes denoting derivatives with respect to the scale factor. PM has negative energy density ($\Omega_{X}<0$) and positive pressure (due to $w_{X}<-1<0$) and, therefore, it induces a decrease of the friction term and an increase of the Poisson term (the last one) in Eq. (4). Both effects are in harmony and therefore bring about, given some fixed initial conditions, a net enhancement of the structure formation processes in the PM bubbles. Notice that this would not occur for ordinary phantom DE nor for quintessence (cf. Fig. 1), for which the friction term gets enhanced and the Poisson term suppressed, i.e. just the opposite situation to PM. This conspicuous impact on LSS is therefore unique to PM since there is no other cosmic fluid capable of matching such an achievement in the EoS diagram of Fig. 1.
@@
Moreover, and this is crucial to understanding how the $H_{0}$ tension can be potentially cured in the PM framework: because the energy of the $X$ entity is negative, $\Omega_{X}<0$ (and non-negligible at high redshift), it enforces a higher value of the expansion rate $H$ in the quintessence stage in order to preserve the angular diameter distance to the last scattering surface, which is essentially fixed from the very precise measurement of $\theta_{*}$ (the angular size of the sound horizon) by Planck (Aghanim et al., 2020) and the standard physics before recombination (Gómez-Valent et al., 2024). This explains on plausible physical grounds why $H_{0}$ is found larger than in the $\Lambda$CDM in our PM scenario. Quantitatively, our value of $H_{0}$ (cf. Table 1) is in full agreement with that of SH0ES ($H_{0}=73.04\pm 1.04$ km/s/Mpc) (Riess et al., 2022) to within $\lesssim 0.25\sigma$. The Hubble tension is therefore virtually washed out, also when it is formulated in terms of the absolute magnitude of SNIa: our fitted value of $M$ and that of SH0ES ($M=19.253\pm 0.027$ mag) differ by only $\sim 0.6\sigma$. At the same time the rate of LSS formation becomes suppressed below $z_{t}\sim 1.5$ during the quintessence-like regime, as explained above, which is also in accordance with the observations. In Table 1 and, more graphically, in Fig. 2, we can see that the amplitude of the power spectrum at linear scales that is preferred by the data is pretty similar in all the models under study. Indeed, in all cases we find values of $\sigma_{12}\sim 0.78$. However, at a finer level of scrutiny, Table 2 in the Appendix tells us that the $w$XCDM is able to describe better the LSS data than the $\Lambda$CDM, what means that the composite DE model is able to produce lower values of $f(z)\sigma_{12}(z)$ and, hence, of $f(z)$ since $\sigma_{12}$ remains stable in the various models. This can again be understood by looking at the equation for the density contrast, whose form at $z<z_{t}$ is identical to Eq. (4), but with the replacements $\Omega_{X}\to\Omega_{Y}$ and $w_{X}\to w_{Y}$,
@@
It is clear from Fig. 2 that the values of the matter density parameter $\Omega_{m}^{0}=\Omega_{m}(a=1)$ in the $w$XCDM are a way smaller than in $\Lambda$CDM. This obviously translates into larger values of $\Omega_{Y}(a)>0$ (higher than $\Omega_{\Lambda}^{0}$ in the $\Lambda$CDM), which makes matter fluctuations to grow less efficiently in our model after the transition (at $z<z_{t}$), a welcome feature that is rightly aligned for potentially solving also the growth tension within the PM picture. As previously mentioned, despite the presence of PM in the past, the value of $\sigma_{12}$ in $w$XCDM is not significantly different ($\sim 0.4\sigma$) from that of $\Lambda$CDM. This is due to the slower increase of matter fluctuations in the late-time universe and the smaller value of the amplitude of primordial fluctuations ($A_{s}$) found in $w$XCDM, which somehow compensate for the presence of PM at $z>z_{t}$ an keep the value of $\sigma_{12}$ stable (cf. again Table 1, and Fig. 4 in the Appendix).
@@
From our numerical analysis we can see that both models $w$XCDM and $\Lambda_{s}$ offer a dramatic reduction of the cosmological tensions. This is apparent from Table 1. One could say that the Hubble and growth tensions disappear in these models. However, there are also important quantitative (and conceptual) differences between them. In terms of the information criteria, we find that the two models $w$XCDM and $\Lambda_{s}$ are very strongly preferred over $\Lambda$CDM. Numerically, however, the $w$XCDM solution provides a significantly better global fit. For this model we find $\Delta$DIC,$\Delta$AIC55greater-than-or-equivalent-toabsent55$\gtrsim$ 55≳ 55, whereas for $\Lambda_{s}$ we obtain $\Delta$DIC,$\Delta$AIC40greater-than-or-equivalent-toabsent40$\gtrsim$ 40≳ 40. The difference of $15$ units is quite substantive and indicates (in the common parlance of the information criteria) that there is very strong evidence that model $w$XCDM is preferred over model $\Lambda_{s}$ (for the same data). Worth noticing is also the fact that our own results for $\Lambda_{s}$ agree with those originally reported in (Akarsu et al., 2023), despite the existing differences in the data sets employed in the two analyses. Apart from our using data on CCH, which were not considered in (Akarsu et al., 2023), we employ data on $f(z)\sigma_{12}(z)$ instead of the weak lensing (WL) data from KiDS. We do not use WL data in our analysis in order not to compromise the computation of the non-linear matter power spectrum, which might not be sufficiently well under control in the models under study. Our fitting results, though, are completely consistent with theirs, leading also to similar values of $S_{8}\sim 0.78$ (as we have checked), the latter being a more natural LSS parameter in the context of WL analyses, but not in our approach.
@@
Perhaps the most remarkable outcome of our analysis is that, in the light of the overall fit quality of our results, which lead to such an outstanding support to the composite DE models from the information criteria, it should be fair to conclude that the standard $\Lambda$CDM model appears to be comparatively ruled out at a very high confidence level. However, once more, we have to warn the reader that, in principle, such a conclusion ensues under the assumption that only BAO 2D have been employed in our analysis (see, however, the additional comments in the Conclusions, i.e. Section 6). In an attempt to further understand the great success of the numerical fits from the composite DE models in Table 1, let us analyze the degree of impact from each data source. The quantitative influence of each data source can be inferred directly from Table 2 in the Appendix, where a detailed breakdown of the different contributions to $\chi^{2}$ is reckoned. The angular/transversal BAO data certainly contributes in a significant way, as could be expected from the analysis of (Gómez-Valent et al., 2024). However, it is by no means the leading contribution, as it is responsible for roughly $\sim 18\%$ of the success. A close inspection of the mentioned table shows that the composite DE models beat actually the $\Lambda$CDM in every single observable, and not only in those driving the tensions. Particularly noticeable is the sizeable impact from such a solid asset of basic data as the CMB and SNIa+SH0ES observations, which are responsible for about $70\%$ of our successful fitting result, whereas the influence from LSS data is less than $10\%$; and, finally, that of CCH is marginal (a few percent). The bulk of the successful fit, therefore, relies on the most fundamental cosmological data, which is remarkable. The $\Lambda$CDM is not only unable to explain the large value of $H_{0}$ measured by SH0ES or to further slow down the growth rate at low redshift, but it also introduces other tensions, e.g. with the value of the (reduced) cosmological matter parameter $\omega_{\rm m}=\omega_{b}+\omega_{\rm dm}$ that is preferred by the Planck data in models which assume standard model physics before the decoupling of the CMB photons, $\omega_{m}=0.142\pm 0.001$ (Aghanim et al., 2020). The latter is fully consistent with the $w$XCDM value ($\omega_{m}=0.142\pm 0.001$), but differs by $2.8\sigma$ with the one inferred in the standard $\Lambda$CDM ($\omega_{m}=0.138\pm 0.001$). All that said, we remark that the BAO data play indeed a momentous role in the $H_{0}$ tension, for despite the fact that the overall fit with composite DE does substantially improve using any of the two variants of BAO, the specific relieve of the $H_{0}$ tension hinges dramatically upon the particular sort (2D or 3D) of BAO that is employed, see Section 6.
@@
Figure 3: Confidence regions in the $w_{Y}-w_{X}$ plane of the $w$XCDM model, and the corresponding one-dimensional posterior distributions. The dotted lines are set at $w_{Y}=-1$ and $w_{X}=-1$. The intersection of the horizontal and vertical lines in the contour plot corresponds to the $\Lambda_{s}$ model, which falls $\gtrsim 3\sigma$ away from the preferred region of the $w$XCDM. See the main text for further comments.
@@
Focusing now on the $w$XCDM parameters, in Fig. 3 we show the constraints obtained in the EoS plane $w_{Y}$-$w_{X}$, which involves two of the characteristic new parameters of the $w$XCDM model (the third one being $z_{t}$, shared with $\Lambda_{s}$). The central value of $w_{X}=-1.16$ falls in the phantom region (in fact, PM region since $\Omega_{X}<0$) but is compatible with $-1$ at $\sim 1\sigma$ c.l. There is, in contrast, a non-negligible ($\sim 3.3\sigma$) preference for a quintessence-like evolution of the DE for the low redshift range nearer to our time ($z<z_{t}$): $w_{Y}=-0.90\pm 0.03$ (see also Fig. 3). The preference that we find for quintessence-like behavior in this region can be contrasted with the situation in the $\Lambda_{s}$, where a rigid $\Lambda$-term is assumed below the transition redshift (also above that redshift, with $\Lambda\to-\Lambda$). Now the fact that our fit with model $w$XCDM proves significantly better than with model $\Lambda_{s}$ (as commented above) suggests that indeed the quintessence option in the low redshift range is strongly preferred. Our result aligns perfectly well with the one obtained from the analysis of the Pantheon+ data in the context of the flat $w$CDM parametrization (Brout et al., 2022), which showed that SNIa data alone lead to an EoS parameter $w=-0.89\pm 0.13$ after marginalizing over $\Omega_{m}^{0}$. This marginalized result, though, is still compatible with a CC ($w=-1$). To explain why in our case is more focused on the quintessence domain, let us note that for $w$XCDM we find a tight constraint on the matter density parameter, strongly favoring the region of small values, $\Omega_{m}^{0}=0.269\pm 0.005$. This is induced by two facts: (i) CMB prefers values of $\omega_{m}\sim 0.142$ in models with standard pre-recombination physics; and (ii) the large values of $H_{0}\sim 73$ km/s/Mpc measured by SH0ES. Combining the aforementioned tight constraint on $\Omega_{m}^{0}$ with the constraints in the $w-\Omega_{m}^{0}$ plane obtained for the $w$CDM from the analysis of the Pantheon+ data alone (cf. the contours in cyan of Fig. 9 of (Brout et al., 2022)) we can break the existing degeneracy in that plane and explain the $\gtrsim 3\sigma$ evidence for quintessence DE that we find within the $w$XCDM for the low redshift range. We point out that, for the conventional $w$CDM parameterization (i.e. without the $X$ component), such low values of $\Omega_{m}^{0}$ are not favored by Planck, since they are in tension with the angular diameter distance to the last scattering surface, see again Fig. 9 of (Brout et al., 2022). Remarkably, this problem can be fully averted in the $w$XCDM model owing precisely to the concurrence of the $X$ component, which can act as PM in the high redshift stretch $z>z_{t}$ and compensate for the decrease of the contribution to the angular diameter distance in the low redshift domain. On the other hand, in the $\Lambda_{s}$ case we also find similar low values of $\Omega_{m}^{0}$. However, to retrieve this model from the $w$XCDM we have to set $w_{Y}$ to $-1$ (rigid CC) [^7] Note that the $\Lambda_{s}$ model can be viewed as just the single point $(w_{Y},w_{X})=(-1,-1)$ in the $w$XCDM parameter space of Fig. 3., so the $\Lambda_{s}$ falls outside the $3\sigma$ region obtained with $w$CDM using only SNIa data (Brout et al., 2022). This issue is also solved in the $w$XCDM and explains the improvement found with respect to the $\Lambda_{s}$ model. The full 5-year data set of high-redshift SNIa from the Dark Energy Survey (Abbott et al., 2024) and the Union3 SNIa compilation (Rubin et al., 2023) will most likely increase the aforesaid signal, since they also find constraints in the region of interest, which read $(\Omega_{m}^{0},w)=(0.264^{+0.074}_{-0.096},-0.80^{+0.14}_{-0.16})$ and $(\Omega_{m}^{0},w)=(0.244^{+0.092}_{-0.128},-0.735^{+0.169}_{-0.191})$, respectively. Hence, it will be of utmost importance to study the impact of these data sets in future analyses. In addition, our fitting results for $w$XCDM are compatible at roughly $\sim 1\sigma$ with those recently reported by DESI using the $w$CDM parametrization (Adame et al., 2024). Although two of the twelve DESI BAO data points (specifically those obtained with Ly-$\alpha$ quasars (QSO) at $z_{\rm eff}=2.33$) lie outside the quintessence region identified in our $w$XCDM model, the bulk of the DESI points lie below $z=1.5$. The exclusion of the Ly-$\alpha$ QSO data points in their fitting analysis might even improve further the compatibility with our findings. Additionally, some recent analyses using DESI data to reconstruct the dark energy density have found hints of negative values at $z>1.5$ (Calderón et al., 2024; Orchard & Cárdenas, 2024). Since the DESI data are derived from anisotropic (3D) BAO, exploring future angular BAO data from DESI (if available) and comparing them with the BAO 3D data would be an interesting avenue for further research.
6 Conclusions
In this paper, we have addressed a possible solution to the cosmological tensions within a simplified version of the old existing $\Lambda$XCDM framework (Grande et al., 2006, 2007, 2009), which was born in the theoretical arena of the RVM – see (Solà Peracaula, 2022; Solà, 2013) and references therein. Our modern approach benefits from the new theoretical developments in the context of the stringy version of the running vacuum model (StRVM) (Mavromatos & Solà Peracaula, 2021b, a). The entity $X$ that is involved in the $\Lambda$XCDM need not be a fundamental field but can display a phantom matter (PM) behaviour, which in the StRVM is predicted to appear in the cosmic transit to the de Sitter phase. The PM regime is, therefore, only transitory and acts as a temporary stage with negative energy and positive pressure. It can be a pure anti-de Sitter (AdS) phase but in general its EoS $w_{X}$ is not necessarily $-1$, although it must satisfy $w_{X}\lesssim-1$. In this work we have contented ourselves with a simplified version of the full $\Lambda$XCDM, which we have called the $w$XCDM. We have utilized a robust data set, which includes the Pantheon+ compilation of SNIa, cosmic chronometers, transverse BAO, large-scale structure data, and the full CMB likelihood from Planck 2018. As warned previously, our exclusive use of transverse (2D) BAO aims at maximally mitigating the model-dependent effects that may be introduced on using anisotropic (3D) BAO, according to some published studies. No similar issues have been identified in BAO 2D and hence we have focused here on using only uncontested BAO data, with the hope that the situation with BAO 3D will be completely clarified in the near future (see, however, the additional comments below).
Using the $w$XCDM and the aforementioned data configuration, we find that the PM phase shows up above a transition redshift $z_{t}\simeq 1.5$ (cf. Table 1). For $z>z_{t}$ the PM behavior is rapidly washed out since its energy density fraction becomes smaller and smaller in the past. And, on the other hand, its effect towards our time ($z<z_{t}$) becomes erased by the progressive appearance of the Chern-Simons condensates, which are responsible for eventually establishing a steady $\Lambda=$const.$>0$ regime (Mavromatos & Solà Peracaula, 2021b). For this reason, in practice, the appearance of a PM phase is confined to a bubble of spacetime, which is left behind. However, in that space and time new structures can emerge, being completely unsuspected in the context of the standard model. The PM bubbles that are met within the stringy RVM formulation originate as a result of a “phantom vacuum” phase, which emerges in the universe when the expansion heads towards the next de Sitter epoch. While the implications of this fact for the early universe were considered previously (Mavromatos & Solà Peracaula, 2021b), here we have extended this phenomenon to the late universe. In both cases the PM bubbles are localized spacetime events. They can only occur as tunneling attempts to establish the phantom vacuum (Mavromatos & Solà Peracaula, 2021b) with positive pressure and negative energy: $p=-\rho>0$. Below the transition redshift, the PM phase ceases to occur and the field $Y$ (viz. the would-be running VED in the full $\Lambda$XCDM (Grande et al., 2006)) takes its turn with a new effective EoS, which is fitted to be of quintessence type ($w_{Y}\gtrsim-1$) at $\sim 3.3\sigma$ c.l. As a matter of fact, this is the effective form of DE that is expected in our most recent past ($z<1.5$) within the $w$XCDM, although from a more fundamental level it could just be running vacuum energy with $\nu_{\rm eff}>0$ in the full $\Lambda$XCDM model, see Eq. (1). The observed DE should thereupon be dynamical, and this important conclusion is in stark contrast with the alternative $\Lambda_{s}$ scheme analyzed in (Akarsu et al., 2023), in which below the transition redshift there is a rigid and positive cosmological term $\Lambda=$const. Actually, no possible time dynamics for the DE is available for the $\Lambda_{s}$, neither above nor below $z_{t}$. On this account, the picture emerging from the $w$XCDM looks more consistent with the first data release by DESI (Adame et al., 2024), which points to dynamical DE. As noted in the Introduction, the fashionable conclusion about potential evidence of time-evolving DE is actually well in consonance with previous phenomenological studies of the RVM from several years ago, which were based on a pretty large set of cosmological data of various sorts. These early detailed studies already anticipated significant signs of dynamical DE at $3\sigma-4\sigma$ c.l. (Solà et al., 2015, 2017; Solà Peracaula et al., 2017, 2018a, 2018b; Gómez-Valent & Solà Peracaula, 2018). The current work aligns well with these studies, which were like a harbinger of the dynamical character of the DE emerging from the analysis of a large body of observational data. But at the same time it leads to further insight into possibly new qualitative properties of the DE, in particular its preference for undergoing a sign flip at a nearby transition redshift, thus adopting a sort of chameleonic nature around that transition point, for the DE appears phantom-like first (viz. “phantom matter”) and subsequently it shows up as effective quintessence near our time (as recently observed by DESI). This combined feature of the DE optimizes to a large extent the quality fit to the overall cosmological data as compared to the $\Lambda$CDM. Now while this positive feature holds true for the two sorts of BAO, it is not so for curing the $H_{0}$ tension (see below).